# Multi-Robot Navigation in Social Mini-Games: Definitions, Taxonomy, and Algorithms

### v3

@@ 3 Algorithms, Taxonomy and Properties for SMGs
We analyze algorithms and system design choices conditioned on an active SMG, i.e., when at least one SMG property from Sec. 2.2 is present (bottleneck contention, crossing flows, persistent mutual occupancy, occlusion, or conflicting objectives). Let denote the active coupling set, that is, the minimal subset of agents whose best responses are mutually coupled during (Sec. 2.2). Coordination, communication, observability, invasiveness, cooperation, and deadlock handling are interpreted on , not on the full set of agents. Throughout, we tie design choices to game-aware outcomes introduced in Sec. 2.3: Social Welfare Gap (SWG), Turn-Taking Fairness (TTF), and Invasiveness Score (IS), reported alongside standard MRN metrics. This section thus specifies how MRN methods become SMG solvers once a trigger activates and a coupling set emerges.
@@ 3.1.1 Multi-Agent Reinforcement Learning (MARL)
MARL extends traditional reinforcement learning to environments with multiple learning agents, operating within the stochastic games framework where the key challenge is that each agent’s learning process is affected by other learning agents. MARL algorithms are categorized as- a) Value-based b) Multi-Agent Deep Q-Networks c) Policy-based. A popular Value-based method is Q-Learning [38], where each agent’s Q-value update rule takes in account of the optimal policy for all agents at equilibrium, governed by an evaluation operator. Multi-Agent Deep Q-Networks (MADQN) [5] extend Deep Q-Networks to multi-agent settings. Value-based methods suffer from scalability issues due to joint action space. On the other hand, Policy-based methods can learn a parametrized policy for each agent . This is the essence of MADDPG [39], where each agent learns its own deterministic or stochastic policy. A fundamental challenge in MARL is solving stochastic games, which can be interpreted as a sequence of normal-form (matrix) games. Often the agents seek Nash equilibrium (NE) solutions where no agent can unilaterally improve their performance by deviating from their policy. The NE can also be directly fed into the Q-learning update by replacing the evaluation operator with a Nash-based operator [40]. However finding and converging to such equilibrium remains computationally challenging. Consequently, one can replace explicit equilibrium computation with learned policies guarded by a reachability certified safety module. Layered Safe MARL [41] trains a policy to minimize multi-agent conflicts and then applies a CBVF [42] reachability based safety filter to resolve the hardest pairwise interactions, yielding safe, efficient navigation in dense traffic and drone swarm hardware tests.
SMG implications (MARL). In active SMGs, MARL naturally models strategic coupling and heterogeneous costs, but requires safety layers (e.g., CBF filters) to avoid catastrophic off-equilibrium play, symmetry-breaking priors or norms to reduce inefficient equilibria, and partial observability handling (occlusions) via belief or prediction modules. With respect to SMGs, MARL can improve fairness and reduce IS.
@@ 3.1.2 Multi-Agent Path Finding (MAPF)
MAPF algorithms solve the problem of computing conflict-free paths for multiple agents navigating a graph. MAPF problems are discrete, centralized, fully cooperative and observable, and entails a global cost function. Two primary optimization criteria dominate MAPF research: sum-of-cost, which minimizes the total steps taken by all agents, and makespan, which minimizes the time for the last agent to reach its goal. MAPF algorithms must balance three fundamental properties that create inherent trade-offs. Optimality ensures minimum global cost, but can come with exponential computational complexity. An example of this is the K-agent A* which scales exponentially with the number of agents. Efficiency prioritizes fast execution, exemplified by Prioritized Planning [43] that assigns sequential movement order to agents, though this approach sacrifices optimality as later agents spend excessive time waiting. Completeness guarantees finding a solution when one exists or correctly reporting failure. Operator Decomposition is an algorithmic approach within the context of MAPF that addresses scalability by breaking the complex multi-agent planning into manageable components called operators. This decomposition allows us to run the A* independently on each operator, thus reducing branching factor. Conflict-based Search [44] employs a two-level approach with a high-level conflict tree construction and a low-level constraint-based path re-computation, enabling optimal solutions while managing computational complexity. The Increasing Tree Cost Search [45] treats MAPG as a lookup problem, maintaining a dictionary of optimal costs and systematically increasing individual agent costs until collision-free solutions emerge. Optimal Reciprocal Collision Avoidance (ORCA-MAPF) [46] bridges discrete MAPF with continuous collision avoidance by combining A*-generated initial paths with velocity obstacles to select a speed for an agent to enable real-time deadlock resolution. Social-MAPF uses auctions to enable the agents to bid dynamically on paths/resources promoting a more balanced distribution, and thus promotes fairness.
@@
SMG implications (MAPF). Classical MAPF is discrete, centralized, fully cooperative and observable. In SMGs, it is best used as a local joint-planning subroutine on : to approximate a socially optimal joint action at bottlenecks/crossings, or to resolve deadlock by injecting priorities. Continuous-time and non-holonomic feasibility is enforced by a short-horizon motion layer. We can achieve strong fairness guarantees when is small and communication is available.
@@ 3.1.3 Optimization
Optimization-based methods form a central paradigm in multi-agent planning, offering a principled way to generate coordinated, efficient, and safe behaviors for teams of robots. These approaches frame the planning problem as the minimization (or maximization) of a cost (or reward) function subject to system dynamics and various constraints, such as collision avoidance, kinematic limits, and task requirements. Optimization-based methods underpin key multi-agent collision avoidance strategies. A key metric fundamentally used for these algorithms is of Velocity Obstacle (VO). VO is the set of velocities to causes an agent to collide with other agents, assuming that other agents maintain their current velocities. Reciprocal Velocity Obstacle (RVO) [47] presents a solution by sharing the responsibility of collision avoidance, through adding a correction velocity to both the agents. However, more the correction velocity increases, more an agent is distracted to the side the more conservative the system becomes. Optimal Reciprocal Collision Avoidance (ORCA) [48] extends the RVO to general n-body collision avoidance problem. ORCA can also be extended to non-holonomic constraints on linear and the angular velocities of the agents too. While VO-based schemes operate in velocity space and can become conservative in cluttered scenes, an alternative is to reason directly in configuration space by decomposing free space into convex sets that integrate naturally with MPC. Free Space Ellipsoid Graphs [49] couple a convex decomposition of the workspace with model-predictive control by reusing the same ellipsoids as graph nodes and state constraints, aligning high-level coordination with low-level safety. To pair global convex partitions with fast local reactions, the control update itself can be formulated as a tractable per-agent convex optimization for reciprocal avoidance under uncertainty. CARP [50] formulates reciprocal multi-robot collision avoidance under asymmetric state uncertainty as a per-agent convex program and extends it to smooth polynomial trajectories amenable to high-rate onboard execution. Beyond per step convex avoidance, coupling intent inference with a runtime safety layer enables exploratory interaction without sacrificing formal guarantees. Shielding-aware dual-control planner couples implicit stochastic MPC for active intent inference with a runtime safety filter, balancing exploration with guaranteed safety during interaction planning [51].
Game theory is the field of mathematical study of strategic interactions and decision-making here each player’s action influences the outcome of all the players in the game. Games are categorized by different criteria. Static games are where all players make a single decision without knowing what decisions other players are making. An example of static game is the prisoner’s dilemma. On the other hand, dynamic games are sequential in nature. Poker is an example of a dynamic game. A zero-sum game refers to one where one player doing better implies that other players will do worse, while a general sum game refer to those where this does not hold. Wang et al. [52] proposed a game-theoretic planning framework where each agent is risk-aware. The approach formulates the multi-agent planning problem as a risk-sensitive dynamic game, where each agent seeks to minimize a risk embedded cost function. The proposed algorithm iteratively approximates the feedback Nash Equilibrium using linearizations of the system dynamics and quadratic approximations of the cost. Iterative linear quadratic games approximate general-sum dynamic games by repeatedly linearizing the dynamics and quadratizing the costs to compute feedback Nash strategies, yielding interactive policies suitable for multi-robot collision avoidance and merging scenarios [53]. A Stackelberg formulation addresses action ordering where branch and play searches over leader follower orders with a mixed-integer planner coupled to trajectory optimization, producing socially efficient, collision free multi-robot plans in crowded interactions [54]. To scale stochastic dynamic games, a partial-belief iLQG variant selectively propagates only the most informative beliefs while preserving equilibrium quality interactions, enabling real time multi-agent planning under uncertainty [55].
@@
SMG implications (Optimization/MPC). Optimization-based planners expose SMG structure via explicit constraints (safety, rules-of-the-road) and multi-objective costs (ego progress versus social costs). On , adding priority or turn-taking constraints yields predictable equilibria with low IS and strong fairness; horizons and solver latency limit scalability. Communication tightens coupling, while no-comm variants rely on shared norms encoded as constraints.
@@ 3.1.4 Others
There are several algorithms that doesn’t fall under the category of the three algorithmic categories discussed so far. These contain algorithms which are based on heuristics, evolutionary algorithms, etc. For example, Zhu et al. [56] introduce a Genetic Algorithm-based Topological Optimisation (GATO) framework for multi-robot logistics in agricultural settings, where the operational environment is represented as a discretized topological map. [57] proposed a Flexible System Design approach for automated warehouses, where a multi-agent simulation modeling based approach is used for motion planning. The approach proposed in [58] is based on coordination space framework, where each robot’s trajectory is represented in a high-dimensional space, and collisions are encoded as obstacles in this space. [59] proposes a hierarchical method with a three-player control system designed for real-world factories. They employ the use of time-expanded graphs for coordination and deadlock prevention. This class of methods are characterized by their adaptability, scalability and are well-suited for domains like warehousing, logistics and large-scale robotic systems. To complement graph-based coordination, we consider data driven controllers that reason directly over continuous state–action spaces under uncertainty. A GP-based decentralized planner learns other agent’s actions and safety envelopes online, enforcing individual and joint safety in continuous spaces without shared policies or centralized control [60]. Coupling this with stable neighbor selection in prediction improves both learning stability and trajectory quality. A smooth attention prior for multi-agent trajectory prediction stabilizes which neighbors an agent attends to over time, improving sample efficiency and accuracy on naturalistic driving datasets [61].
SMG implications (Heuristics/Hybrids). Local reactive methods (e.g., ORCA variants) are effective for pairwise interactions but can stall in symmetric SMGs (head-on hallway, doorway) without symmetry-breaking or priority rules. Hybrids (e.g., ORCA+local-MAPF) treat SMG activations as events that trigger a brief centralized or negotiated joint solve on to restore liveness and fairness.
3.2 Taxonomy of SMG Design Choices
@@ 3.2.1 Coordination
In SMGs, coordination among agents is crucial. Coordination refers to how the agents make decisions and take actions within their environment. It includes the mechanisms through which the motion-planner decisions are made, shared, and implemented. The first type of coordination is Centralized which involves a single central authority making decisions for all agents [62, 63]. In this setup, individual agents follow the commands from the central entity, without generating their own controls. This class of planners ensure that all agents work together smoothly and efficiently, as they all follow the same strategy. This kind of coordination is used in methods like Right Hand Rule and Auctions [64], where the central entity generates the command for all the other agents for the safe navigation making sure each agent reaches their goal.
@@
Typical centralized multi-robot pathfinding (MAPF) methods also follow this coordination style. Algorithms like Conflict-Based Search (CBS) [44] and its suboptimal variant Enhanced CBS (ECBS) [65] are widely used, where a central planner coordinates all agents’ paths by resolving conflicts globally. Similarly, methods like M* [66] and ICTS (Increasing Cost Tree Search) [45] compute joint plans that guarantee collision-free paths for all agents by centrally searching the combined state space. These approaches ensure global consistency and optimality (or bounded suboptimality), but their complexity grows rapidly with the number of agents. Therefore, centralized MAPF methods are highly effective in structured environments with moderate agent counts, offering strong coordination guarantees but requiring powerful centralized computation.
@@
Beyond classical centralized planning approaches, recent research has explored Distributed methods that consist of centralized training of agents using reinforcement learning (RL) with communication protocols learned or designed to support coordination [67, 68]. In this setup, agents are trained together using a shared reward signal and centralized critic, but may execute actions independently with limited or explicit communication. Methods such as those developed by Prorok et al. [69, 70] focus on learning decentralized policies that embed communication for distributed coordination. These approaches differ from traditional MAPF methods by allowing agents to dynamically adapt to new scenarios without explicit replanning. By combining centralized optimization during training with communication-aware decentralized execution, such RL-based methods achieve flexible, robust, and scalable multi-robot coordination in complex environments where centralized control at runtime may not be feasible.
@@
The last type of coordination is Decentralized which allows each agent to make its own decisions independently. Instead of relying on a central authority, each agent acts based on its own observations, rewards, and goals. This approach promotes independence and flexibility, enabling agents to adapt to their surroundings and other agents dynamically [71, 72]. Although it may seem less organized, with the right algorithms and planners, decentralized coordination can generate effective and efficient controls to develop a safe, collision-free, and deadlock-free trajectory [73, 36, 74, 75, 76]. Examples of this approach include IMPC-DR [77] and ORCA-MAPF [46], where agents independently navigate in their environment while avoiding collisions and deadlocks. Figures 4 and 5 list the key papers for the Centralized and Decentralized communication categories. Note that the color legend signifies the categorization based on Algorithmic/Methodological paradigm discussed in Sec. 3.1, where we use different colors for different approaches- multi-robot Reinforcement Learning (cyan), Optimization (violet), multi-robot Path Finding (green), and Others (Orange).

### v4

@@ 3 Algorithms, Taxonomy and Properties for SMGs
We analyze algorithms and system design choices conditioned on an active SMG, i.e., when at least one SMG property from Sec. 2.4 is present (bottleneck contention, crossing flows, persistent mutual occupancy, occlusion, or conflicting objectives). Let denote the active coupling set, that is, the minimal subset of agents whose best responses are mutually coupled during (Sec. 2.4). At any time , agent interactions may induce multiple, possibly overlapping regions of strategic coupling. To formalize this, we define an interaction graph , where the node set corresponds to agents, and an undirected edge exists if the best-response policy of agent depends nontrivially on the actions or state of agent over a nonzero time interval containing . We define the coupling sets as the connected components of the interaction graph . Each represents a maximal subset of agents whose decision-making processes are mutually coupled at time . When the context is clear, we use to denote a generic coupling set.
This formulation naturally allows for multiple coupling sets to exist simultaneously. Moreover, coupling sets may overlap across time as interaction contexts evolve. For example, if agents are coupled while traversing doorway A and agents are later coupled while traversing doorway B, then distinct coupling sets and may exist at different times. If the planning or learning algorithm reasons jointly over all interacting agents, the union may be treated as a single coupled group over the corresponding time interval. Throughout this paper, our taxonomy, metrics, and SMG analyses are applied independently to each active coupling set . This enables disjoint social mini-games to be identified, analyzed, and evaluated without requiring global coupling across all agents. Throughout, we tie design choices to game-aware outcomes introduced in Sec. 2.5: Social Welfare Gap (SWG), Turn-Taking Fairness (TTF), and Invasiveness Score (IS), reported alongside standard MRN metrics. This section thus specifies how MRN methods become SMG solvers once a trigger activates and a coupling set emerges.
@@ 3.1.1 Multi-Agent Reinforcement Learning (MARL)
MARL extends single-agent reinforcement learning to environments with multiple learning agents, operating within the stochastic games framework where the key challenge is that each agent’s learning process is affected by other learning agents (also known as the well-known non-stationarity problem). A MARL formulation models the system using the POSG tuple defined in Section 2.1. In MARL, the primary objective of each agent is to minimize an expected cost function: using standard RL techniques, where denote the joint policies and actions of all agents except .
An SMG arises when there exists non-stationarity in a MARL approach. Non-stationarity refers to the scenario where the RL policy of an agent fails to converge because the agent’s environment keeps changing due to varying actions of other agents. That is, when the reward or transition function induces strong interdependence among a subset of agents such that it leads to non-stationarity, then it can be shown that those agents are the ones that typically belong to , and therefore institute a SMG. In particular, the local state relevant to the SMG consists of and environmental features defining the bottleneck (e.g., doorway geometry), the actions are velocity or acceleration commands, and the rewards couple agents through collision penalties, shared resource contention, or delay penalties. Thus, in MARL, an SMG corresponds to a localized stochastic subgame restricted to agents in . In active SMGs, MARL naturally models strategic coupling and heterogeneous costs, but requires safety layers (e.g., CBF filters) to avoid catastrophic off-equilibrium play, symmetry-breaking priors or norms to reduce inefficient equilibria, and partial observability handling (occlusions) via belief or prediction modules. With respect to SMGs, MARL can improve fairness and reduce IS.
MARL algorithms are categorized as- a) Value-based b) Multi-Agent Deep Q-Networks c) Policy-based. A popular Value-based method is Q-Learning [tampuu2017multiagent], where each agent’s Q-value update rule takes in account of the optimal policy for all agents at equilibrium, governed by an evaluation operator. Multi-Agent Deep Q-Networks (MADQN) [orr2023multi] extend Deep Q-Networks to multi-agent settings. Value-based methods suffer from scalability issues due to joint action space. On the other hand, Policy-based methods can learn a parametrized policy for each agent . This is the essence of MADDPG [lowe2017multi], where each agent learns its own deterministic or stochastic policy. A fundamental challenge in MARL is solving stochastic games, which can be interpreted as a sequence of normal-form (matrix) games. Often the agents seek Nash equilibrium (NE) solutions where no agent can unilaterally improve their performance by deviating from their policy. The NE can also be directly fed into the Q-learning update by replacing the evaluation operator with a Nash-based operator [hu2003nash]. However finding and converging to such equilibrium remains computationally challenging. Consequently, one can replace explicit equilibrium computation with learned policies guarded by a reachability certified safety module. Layered Safe MARL [choi2025resolving] trains a policy to minimize multi-agent conflicts and then applies a CBVF [choi2021robust] reachability based safety filter to resolve the hardest pairwise interactions, yielding safe, efficient navigation in dense traffic and drone swarm hardware tests.
@@ 3.1.2 Multi-Agent Path Finding (MAPF)
MAPF algorithms solve the problem of computing conflict-free paths for multiple agents navigating a graph. The environment is modeled as a graph . Each agent has an initial vertex and a goal vertex . At each discrete time step, an agent may move to an adjacent vertex or wait. A solution consists of collision-free paths such that and . Typical objectives include sum of costs (total path length) and Makespan (maximum arrival time).
@@
An SMG appears when multiple agents’ shortest paths traverse a shared low-capacity region (e.g., a single-width corridor node or edge). In this case, the state is the discrete occupancy configuration, actions are graph transitions, and resource contention manifests as vertex or edge conflicts. The coupling set consists of agents whose shortest paths intersect at shared vertices or edges within overlapping time intervals. The SMG is the local discrete coordination problem required to resolve those conflicts. Classical MAPF is discrete, centralized, fully cooperative and observable. In SMGs, it is best used as a local joint-planning subroutine on : to approximate a socially optimal joint action at bottlenecks/crossings, or to resolve deadlock by injecting priorities. Continuous-time and non-holonomic feasibility is enforced by a short-horizon motion layer. We can achieve strong fairness guarantees when is small and communication is available.
MAPF problems are discrete, centralized, fully cooperative and observable, and entails a global cost function. Two primary optimization criteria dominate MAPF research: sum-of-cost, which minimizes the total steps taken by all agents, and makespan, which minimizes the time for the last agent to reach its goal. MAPF algorithms must balance three fundamental properties that create inherent trade-offs. Optimality ensures minimum global cost, but can come with exponential computational complexity. An example of this is the K-agent A* which scales exponentially with the number of agents. Efficiency prioritizes fast execution, exemplified by Prioritized Planning [silver2005cooperative] that assigns sequential movement order to agents, though this approach sacrifices optimality as later agents spend excessive time waiting. Completeness guarantees finding a solution when one exists or correctly reporting failure. Operator Decomposition is an algorithmic approach within the context of MAPF that addresses scalability by breaking the complex multi-agent planning into manageable components called operators. This decomposition allows us to run the A* independently on each operator, thus reducing branching factor. Conflict-based Search [sharon2015conflict] employs a two-level approach with a high-level conflict tree construction and a low-level constraint-based path re-computation, enabling optimal solutions while managing computational complexity. The Increasing Tree Cost Search [sharon2013icts] treats MAPG as a lookup problem, maintaining a dictionary of optimal costs and systematically increasing individual agent costs until collision-free solutions emerge. Optimal Reciprocal Collision Avoidance (ORCA-MAPF) [dergachev2021distributed] bridges discrete MAPF with continuous collision avoidance by combining A*-generated initial paths with velocity obstacles to select a speed for an agent to enable real-time deadlock resolution. Social-MAPF uses auctions to enable the agents to bid dynamically on paths/resources promoting a more balanced distribution, and thus promotes fairness.
@@ 3.1.3 Optimization
Optimization-based approaches formulate navigation as a constrained optimal control problem. For each agent , one solves:
where is the terminal cost. A solution to optimal navigation in social mini-games is a (finite) sequence of state-input pairs . In centralized formulations, the optimization is performed jointly over . In decentralized formulations, collision avoidance appears as constraints coupling agents.
@@
An SMG corresponds to a localized constrained subproblem over agents in , where the decision variables are trajectories or control sequences over a horizon. Constraints encode collision avoidance and possibly resource capacity (e.g., at most one agent inside a doorway region at a time). The coupling set is determined by agents whose constraints are mutually active. Thus, in optimization-based methods, an SMG is characterized by the activation of shared inequality constraints among a subset of agents over a finite time interval. Optimization-based planners expose SMG structure via explicit constraints (safety, rules-of-the-road) and multi-objective costs (ego progress versus social costs). On , adding priority or turn-taking constraints yields predictable equilibria with low IS and strong fairness; horizons and solver latency limit scalability. Communication tightens coupling, while no-comm variants rely on shared norms encoded as constraints.
Optimization-based approaches frame the planning problem as the minimization (or maximization) of a cost (or reward) function subject to system dynamics and various constraints, such as collision avoidance, kinematic limits, and task requirements. Optimization-based methods underpin key multi-agent collision avoidance strategies. A key metric fundamentally used for these algorithms is of Velocity Obstacle (VO). VO is the set of velocities to causes an agent to collide with other agents, assuming that other agents maintain their current velocities. Reciprocal Velocity Obstacle (RVO) [berg2008reciprocal] presents a solution by sharing the responsibility of collision avoidance, through adding a correction velocity to both the agents. However, more the correction velocity increases, more an agent is distracted to the side the more conservative the system becomes. Optimal Reciprocal Collision Avoidance (ORCA) [nh-orca] extends the RVO to general n-body collision avoidance problem. ORCA can also be extended to non-holonomic constraints on linear and the angular velocities of the agents too. While VO-based schemes operate in velocity space and can become conservative in cluttered scenes, an alternative is to reason directly in configuration space by decomposing free space into convex sets that integrate naturally with MPC. Free Space Ellipsoid Graphs [ray2022free] couple a convex decomposition of the workspace with model-predictive control by reusing the same ellipsoids as graph nodes and state constraints, aligning high-level coordination with low-level safety. To pair global convex partitions with fast local reactions, the control update itself can be formulated as a tractable per-agent convex optimization for reciprocal avoidance under uncertainty. CARP [shah2021reciprocal] formulates reciprocal multi-robot collision avoidance under asymmetric state uncertainty as a per-agent convex program and extends it to smooth polynomial trajectories amenable to high-rate onboard execution. Beyond per step convex avoidance, coupling intent inference with a runtime safety layer enables exploratory interaction without sacrificing formal guarantees. Shielding-aware dual-control planner couples implicit stochastic MPC for active intent inference with a runtime safety filter, balancing exploration with guaranteed safety during interaction planning [hu2024active].
Game theory is the field of mathematical study of strategic interactions and decision-making here each player’s action influences the outcome of all the players in the game. Games are categorized by different criteria. Static games are where all players make a single decision without knowing what decisions other players are making. An example of static game is the prisoner’s dilemma. On the other hand, dynamic games are sequential in nature. Poker is an example of a dynamic game. A zero-sum game refers to one where one player doing better implies that other players will do worse, while a general sum game refer to those where this does not hold. Wang et al. [wang2020game] proposed a game-theoretic planning framework where each agent is risk-aware. The approach formulates the multi-agent planning problem as a risk-sensitive dynamic game, where each agent seeks to minimize a risk embedded cost function. The proposed algorithm iteratively approximates the feedback Nash Equilibrium using linearizations of the system dynamics and quadratic approximations of the cost. Iterative linear quadratic games approximate general-sum dynamic games by repeatedly linearizing the dynamics and quadratizing the costs to compute feedback Nash strategies, yielding interactive policies suitable for multi-robot collision avoidance and merging scenarios [fridovich2020efficient]. A Stackelberg formulation addresses action ordering where branch and play searches over leader follower orders with a mixed-integer planner coupled to trajectory optimization, producing socially efficient, collision free multi-robot plans in crowded interactions [hu2024plays]. To scale stochastic dynamic games, a partial-belief iLQG variant selectively propagates only the most informative beliefs while preserving equilibrium quality interactions, enabling real time multi-agent planning under uncertainty [vakil2024partial].
@@ 3.1.4 Others
There are several algorithms that do not fall under the category of the three algorithmic categories discussed so far. These contain algorithms which are based on heuristics, evolutionary algorithms, etc. For example, Zhu et al. [das2023autonomous] introduce a Genetic Algorithm-based Topological Optimisation (GATO) framework for multi-robot logistics in agricultural settings, where the operational environment is represented as a discretized topological map. [turhanlar2021deadlock] proposed a Flexible System Design approach for automated warehouses, where a multi-agent simulation modeling based approach is used for motion planning. The approach proposed in [coskun2021deadlock] is based on coordination space framework, where each robot’s trajectory is represented in a high-dimensional space, and collisions are encoded as obstacles in this space. [pratissoli2023hierarchical] proposes a hierarchical method with a three-player control system designed for real-world factories. They employ the use of time-expanded graphs for coordination and deadlock prevention. This class of methods are characterized by their adaptability, scalability and are well-suited for domains like warehousing, logistics and large-scale robotic systems. To complement graph-based coordination, we consider data driven controllers that reason directly over continuous state–action spaces under uncertainty. A GP-based decentralized planner learns other agent’s actions and safety envelopes online, enforcing individual and joint safety in continuous spaces without shared policies or centralized control [zhu2020multi]. Coupling this with stable neighbor selection in prediction improves both learning stability and trajectory quality. A smooth attention prior for multi-agent trajectory prediction stabilizes which neighbors an agent attends to over time, improving sample efficiency and accuracy on naturalistic driving datasets [cao2022leveraging].
Local reactive methods (e.g., ORCA variants) are effective for pairwise interactions but can stall in symmetric SMGs (head-on hallway, doorway) without symmetry-breaking or priority rules. Hybrids (e.g., ORCA+local-MAPF) treat SMG activations as events that trigger a brief centralized or negotiated joint solve on to restore liveness and fairness.
3.2 Taxonomy of SMG Solvers
@@ 3.2.1 Coordination
In SMGs, coordination among agents is crucial. Coordination refers to how the agents make decisions and take actions within their environment. It includes the mechanisms through which the motion-planner decisions are made, shared, and implemented. The first type of coordination is Centralized which involves a single central authority making decisions for all agents [saha2006multi, li2017centralized]. In this setup, individual agents follow the commands from the central entity, without generating their own controls. This class of planners ensure that all agents work together smoothly and efficiently, as they all follow the same strategy. This kind of coordination is used in methods like Right Hand Rule and Auctions [wang2017safety], where the central entity generates the command for all the other agents for the safe navigation making sure each agent reaches their goal.
@@
Typical centralized multi-robot pathfinding (MAPF) methods also follow this coordination style. Algorithms like Conflict-Based Search (CBS) [sharon2015conflict] and its suboptimal variant Enhanced CBS (ECBS) [barer2014suboptimal] are widely used, where a central planner coordinates all agents’ paths by resolving conflicts globally. Similarly, methods like M* [wagner2011mstar] and ICTS (Increasing Cost Tree Search) [sharon2013icts] compute joint plans that guarantee collision-free paths for all agents by centrally searching the combined state space. These approaches ensure global consistency and optimality (or bounded suboptimality), but their complexity grows rapidly with the number of agents. Therefore, centralized MAPF methods are highly effective in structured environments with moderate agent counts, offering strong coordination guarantees but requiring powerful centralized computation.
@@
Beyond classical centralized planning approaches, recent research has explored Distributed methods that consist of centralized training of agents using reinforcement learning (RL) with communication protocols learned or designed to support coordination [das2019tarmac, sukhbaatar2016commnet]. In this setup, agents are trained together using a shared reward signal and centralized critic, but may execute actions independently with limited or explicit communication. Methods such as those developed by Prorok et al. [cao2021multi, wang2020multi] focus on learning decentralized policies that embed communication for distributed coordination. These approaches differ from traditional MAPF methods by allowing agents to dynamically adapt to new scenarios without explicit replanning. By combining centralized optimization during training with communication-aware decentralized execution, such RL-based methods achieve flexible, robust, and scalable multi-robot coordination in complex environments where centralized control at runtime may not be feasible.
@@
The last type of coordination is Decentralized which allows each agent to make its own decisions independently. Instead of relying on a central authority, each agent acts based on its own observations, rewards, and goals. This approach promotes independence and flexibility, enabling agents to adapt to their surroundings and other agents dynamically [azarm1997conflict, demesure2017decentralized]. Although it may seem less organized, with the right algorithms and planners, decentralized coordination can generate effective and efficient controls to develop a safe, collision-free, and deadlock-free trajectory [das2019sparcas, long2018towards, desaraju2012decentralized, maoudj2023improved, toumieh2022decentralized]. Examples of this approach include IMPC-DR [chen2023multi] and ORCA-MAPF [dergachev2021distributed], where agents independently navigate in their environment while avoiding collisions and deadlocks. Figures 6 and 7 list the key papers for the Centralized and Decentralized communication categories. Note that the color legend signifies the categorization based on Algorithmic/Methodological paradigm discussed in Sec. 3.1, where we use different colors for different approaches- multi-robot Reinforcement Learning (cyan), Optimization (violet), multi-robot Path Finding (green), and Others (Orange).